OpenStudy (anonymous):

**Will Medal** What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y = ? x + ?

OpenStudy (anonymous):

I need help ASAP plzzz!!!!!!!!!!!!!!!!!!

zepdrix (zepdrix):

Our slope-intercept form of a line is given by:\[\Large\rm y=mx+b\]Do you know which letter corresponds to the slope?

zepdrix (zepdrix):

Woops I clicked the wrong question >.< lol Sorry sorry.

zepdrix (zepdrix):

Anyway, y=mx+b. m is our slope.

OpenStudy (anonymous):

Ok

zepdrix (zepdrix):

Giving us,\[\Large\rm y=2x+b\]

zepdrix (zepdrix):

To find the y-intercept, \(\Large\rm b\), we need to plug in the point that they gave us,\[\Large\rm 4=2(3)+b\]And then solve for b. Do you understand how I plugged that coordinate in? Solve for b! What do you get? :)

OpenStudy (anonymous):

So I would multiply 2 and which gives me 6?

OpenStudy (anonymous):

2 and 3 my bad

zepdrix (zepdrix):

mhm, then we need to isolate the b somehow. Hmm, the 6 is being `added` to the b. To isolate the b, we want to perform the inverse to move the 6 to the other side. Inverse of addition? :o

OpenStudy (anonymous):

Um not sure how to do that

zepdrix (zepdrix):

Subtraction is the opposite of addition, yes? So we'll subtract 6 from each side to get the b alone.

OpenStudy (anonymous):

Ok which comes out to 6b or b6

zepdrix (zepdrix):

6b? Oh boy :( Not sure where that's coming from.